Consider the two regression lines 3x+2y =26 and 6x+y=31,(a)Find the mean value and correlation conefficient between x and y.(b) if the varianceof y is 4,Find the S.D .of x


1
Expert's answer
2022-01-04T15:28:02-0500

"3x+2y=26"...i

"6x+y=31" ....ii balancing both y by multiplying by 2

"3x+2y=26"

"12x+2y=62"

Deducting the two equations

"-9x=-36"

"x=4"

replacing the value of x to get y

"6(4)+y=31"

"y=31-24"

"y=7"

Since the point of intersection of the two correlation lines is x,y then the point is (4,7)

Then the mean of x is 4 and y is 7.

Assuming 3x+2y-26=0 is the regression equation.

2y=-3x+26

y=-3/2x+13

Hence the correlation coefficient of y to x is -3/2

Hence the correlation coeffient of x to y is -1/6